OpenStudy (anonymous):

If you have a new deck of shuffled cards. What is the probability you will choose a black card? (52 cards, no jokers)

OpenStudy (anonymous):

1 out of 2

OpenStudy (anonymous):

thanks c:

jimthompson5910 (jim_thompson5910):

hopefully you see why it's 1/2, if not then there are 26 black cards (13 spade + 13 club) out of 52 26/52 = (26*1)/(26*2) = 1/2 So 26/52 = 1/2 which explains why the answer is 1/2 (or 0.5 or 50%)